# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ D.5 Result for Super Mario
Llama-3.2-1B 18.7$\pm$8.6 12
Llama-3.2-3B 31.8$\pm$10.1 4
Qwen-2.5-3B 23.4$\pm$14.1 11
Qwen-2.5-7B 27.2$\pm$9.6 9
Minitron-4B 24.4$\pm$6.0 10
Minitron-8B 31.3$\pm$12.8 6
GPT-4o-mini 28.8$\pm$8.8 7
GPT-4o 34.1$\pm$14.2 3
o3-mini 34.9$\pm$14.6 2
Gemini-2.5-pro 38.0$\pm$14.6 1
Claude-3.7 31.7$\pm$8.2 5
Deepseek-R1 28.7$\pm$13.2 8
@@
Llama-3B Zero-shot 21.2$\pm$8.2 8
Reflection 32.4$\pm$8.6 2
Planning 27.0$\pm$8.4 7
Ref-Plan 31.8$\pm$10.1 4
GPT-4o Zero-shot 29.6$\pm$9.2 5
Reflection 32.3$\pm$15.5 3
Planning 29.4$\pm$12.5 6
Ref-Plan 34.1$\pm$14.2 1
@@
GPT-4o Text 34.1$\pm$14.1 3
Image 27.1$\pm$13.7 6.5
Both 27.1$\pm$10.2 6.5
Gemini Text 38.0$\pm$13.4 2
Image 28.5$\pm$10.7 5
Both 40.9$\pm$9.6 1
Claude Text 28.7$\pm$13.2 4
Image 25.6$\pm$6.4 8
Both 22.6$\pm$6.3 9
@@ E.1 Game Description for Ace Attorney
Environment. Ace Attorney [55] is a courtroom adventure game where players act as defense attorneys, gather evidence, and cross-examine witnesses. We target the first episode of Phoenix Wright: Ace Attorney Trilogy on Steam (see Figure 12). We define four subtasks: one three-question multiple-choice quiz where the player selects the correct answer, and three cross-examination tasks where the player presses witnesses for more details or presents evidence to expose contradictions. We use Harmony [78] with a BepInEx plugin [79] to hook the game’s source code at launch, capturing states such as dialogue text, arrow-button visibility, keyboard inputs, and Court Record entries. The hooks save states as .txt or .json files and monitor a command .txt file for inputs, injecting them into the game in real time.
Observation-to-Text Conversion. All states remain in text form and require only minimal post-processing. We map speaker indices to character names (e.g., index ‘2’ $\rightarrow$ ‘Phoenix Wright’) using a predefined mapping and replace original names with arbitrary aliases to prevent contamination (e.g., ‘Phoenix Wright’ $\rightarrow$ ‘Alias’). We also convert Court Records and multiple-choice candidate options—originally stored in .json—into continuous descriptive text when they exist.
@@ E.3 Evaluation Metric for Ace Attorney
Each subtask begins and ends at fixed points, with screenshots at both start and end and the preceding conversation history provided, and yields a reward $r_{i}$ and a step count $t_{i}$. In the multiple-choice task (MC), $r_{i}$ is the number of correct answers out of three, while in cross-examination tasks CE1, CE2, and CE3, $r_{i}$ is 1 for a pass and 0 for a fail; any failure incurs a maximum step count of 50. We normalize each reward and step count against fixed benchmarks $\bar{r}_{i}$ and $\bar{t}_{i}$, namely $(3,25)$ for MC, $(1,11)$ for CE1, $(1,3)$ for CE2, and $(1,4)$ for CE3, by computing
@@ E.5 Result for Ace Attorney
Llama-3.2-1B 1.3$\pm$2.2 12
Llama-3.2-3B 4.6$\pm$1.3 11
Qwen-2.5-3B 20.0$\pm$17.4 9
Qwen-2.5-7B 9.3$\pm$0.2 10
Minitron-4B 35.7$\pm$4.5 6
Minitron-8B 29.9$\pm$3.6 7
GPT-4o-mini 28.4$\pm$2.8 8
GPT-4o 85.3$\pm$1.5 2
o3-mini 91.7$\pm$1.5 1
Gemini-2.5-pro 55.7$\pm$3.4 5
Claude-3.7 81.9$\pm$1.6 4
Deepseek-R1 83.3$\pm$1.5 3
@@
Llama-3B Zero-shot 5.7$\pm$3.2 5
Reflection 4.6$\pm$1.3 6.5
Planning 4.6$\pm$1.3 6.5
Ref-Plan 3.8$\pm$0.0 8
GPT-4o Zero-shot 49.9$\pm$1.3 4
Reflection 85.3$\pm$1.5 1
Planning 52.7$\pm$0.5 3
Ref-Plan 52.8$\pm$0.5 2
@@
GPT-4o Text 85.3$\pm$1.5 1
Both 53.5$\pm$1.7 5
Gemini Text 55.7$\pm$3.4 4
Both 52.6$\pm$0.8 6
Claude Text 81.9$\pm$1.6 2
Both 71.3$\pm$17.3 3
@@ F.1 Game Description for Her Story
Environment. Her Story [56] is an interactive adventure game where players explore police interview clips to uncover a hidden truth. The player begins the game by accessing an old desktop interface, where a program called L.O.G.I.C. Database is open. The player can enter keywords into the database to retrieve up to five video clips whose transcripts contain the searched word. By watching these clips and gathering clues, the player repeatedly formulates new queries to reconstruct the underlying story. To interface the game with our code, we use Harmony [78] together with Unity Doorstop [80] to log internal game states to a .txt file. Specifically, each line in the file is a JSON object representing a snapshot of the game’s state at a given moment. Each object contains an event type and its associated metadata. The following examples illustrate key elements of the logged game state. For clarity, we omit auxiliary metadata that are present in the actual logs but are only used for debugging or UI-related purposes, such as video IDs, screen resolution, and UI element positions.
  • Load title screen: {"status": "title"}
  • Load L.O.G.I.C. Database: {"status": "start_game"}
  • Query: {"status": "query", "keyword": keyword}
  • Get query result: {"status": "query_result", "num_total": number of clips containing the keyword, "num_visible": number of clips shown, "video": list of {"new": 1 if not viewed, otherwise 0, "session": recording date, "outfit": visual description of the thumbnail}}
  • Open video panel: {"status": "open_detail"}
  • Close video panel: {"status": "close_detail"}
  • Play video: {"status": "play_video", "script": transcript of the video}
  • Close video: {"status": "close_video"}
@@ F.3 Evaluation Metric for Her Story
To uncover the truth behind the case, the player actively explores the video archive by issuing queries and watching clips. We evaluate this behavior by counting how many unique clips the player has viewed from the full archive of 272 clips. This aligns with the original game design, where unlocking achievements and reaching the ending depend on the number of clips watched. Specifically, we define the score as:
Score=(Number of distinct video clips $\text{Score}=\left(\frac{\text{Number of distinct video clips viewed}}{272}\right)\times 100.$
@@ F.5 Result for Her Story
Table 22 summarizes the gameplay scores of different LLMs on Her Story. Commercial LLMs outperform open-source models in this task. Gemini-2.5-pro achieves the highest score of 67.3, followed by Deepseek-R1 (66.9), o3-mini (66.3), GPT-4o (64.0), Claude-3.7-sonnet (62.6), and GPT-4o-mini (21.1). The low score of GPT-4o-mini is due to its repeated search of keywords such as alibi, evidence, and witness, even after those queries have already been issued. This redundancy reduces its ability to discover new clips. Among open-source LLMs, all models score below 10. Most exhibited similar failure patterns: they fail to play unseen clips, repeatedly issue the same keywords, or search using entire sentences (e.g., for the name "Simon", for keywords related to Luna’s) rather than meaningful single words.
@@
Llama-3.2-1B 2.1$\pm$1.2 11
Llama-3.2-3B 4.2$\pm$1.1 10
Qwen-2.5-3B 1.2$\pm$1.1 12
Qwen-2.5-7B 8.5$\pm$1.9 7
Minitron-4B 4.5$\pm$2.2 9
Minitron-8B 8.2$\pm$1.8 8
GPT-4o-mini 21.1$\pm$5.5 6
GPT-4o 64.0$\pm$5.1 4
o3-mini 66.3$\pm$3.6 3
Gemini-2.5-pro 67.3$\pm$3.3 1
Claude-3.7 62.6$\pm$2.6 5
Deepseek-R1 66.9$\pm$3.9 2
@@
Llama-3B Zero-shot 4.2$\pm$1.1 8
Reflection 4.4$\pm$1.3 7
Planning 5.1$\pm$1.0 6
Ref-Plan 5.4$\pm$0.4 5
GPT-4o Zero-shot 64.0$\pm$5.1 1
Reflection 61.3$\pm$0.4 3
Planning 59.2$\pm$5.7 4
Ref-Plan 61.8$\pm$4.9 2
@@
GPT-4o Text 64.0$\pm$5.1 3
Both 40.4$\pm$29.4 6
Gemini Text 67.3$\pm$3.3 1
Both 64.7$\pm$2.4 2
Claude Text 62.6$\pm$2.6 5
Both 63.4$\pm$3.1 4
@@ G.1 Game Description for Pokémon Red
Environment. Pokémon Red [57] is a role-playing game where the player navigates the Kanto region to catch and train creatures called Pokémon, battle other trainers, and ultimately defeat the Elite Four and the Champion. The player explores various environments, including towns, routes, caves, and buildings, encountering wild Pokémon and other characters. The gameplay loop involves exploring these areas, engaging in turn-based battles with Pokémon, and managing a team of up to six Pokémon. For implementation, we utilize the PyBoy [81] emulator to run the game. Specifically, our evaluation focuses on a segment where the player starts in Pallet Town and progresses towards Viridian City, encountering wild Pokémon and trainers. The game screen resolution is 160$\times$144 pixels.
@@ G.3 Evaluation Metric for Pokémon Red
Score=(Number of flags $\text{Score}=\big{(}\frac{\text{Number of flags achieved}}{12}\big{)}\times 100.$
@@ G.5 Result for Pokémon Red
Llama-3.2-1B 0.0$\pm$0.0 8.5
Llama-3.2-3B 0.0$\pm$0.0 8.5
Qwen-2.5-3B 0.0$\pm$0.0 8.5
Qwen-2.5-7B 0.0$\pm$0.0 8.5
Minitron-4B 0.0$\pm$0.0 8.5
Minitron-8B 0.0$\pm$0.0 8.5
GPT-4o-mini 0.0$\pm$0.0 8.5
GPT-4o 38.9$\pm$9.6 4
o3-mini 0.0$\pm$0.0 8.5
Gemini-2.5-pro 83.3$\pm$0.0 1
Claude-3.7 63.9$\pm$9.2 3
Deepseek-R1 75.0$\pm$0.0 2
@@
Llama-3B Zero-shot 0.0$\pm$0.0 6.5
Reflection 0.0$\pm$0.0 6.5
Planning 0.0$\pm$0.0 6.5
Ref-Plan 0.0$\pm$0.0 6.5
GPT-4o Zero-shot 33.3$\pm$0.0 3.5
Reflection 36.1$\pm$4.8 2
Planning 33.3$\pm$0.0 3.5
Ref-Plan 38.9$\pm$9.6 1
@@
GPT-4o Text 38.9$\pm$9.6 6
Both 41.7$\pm$8.3 5
Gemini Text 83.3$\pm$0.0 1.5
Both 83.3$\pm$0.0 1.5
Claude Text 63.9$\pm$19.2 4
Both 72.2$\pm$4.8 3
@@ H.1 Game Description for Darkest Dungeon
Environment. Darkest Dungeon is a turn-based roguelike role-playing game where the player manages a roster of heroes as they explore procedurally generated dungeons filled with monsters, traps, and treasures  [58]. Each hero has unique abilities and a stress level that influences their behavior during combat and exploration. Stress accumulates through continued exploration and battle, and heroes who reach high stress thresholds may develop afflictions that hinder or occasionally enhance their performance. The game emphasizes tactical positioning, turn-order strategy, and long-term roster management. Elements of randomness, such as attack accuracy, critical hits, and affliction outcomes, introduce uncertainty and require players to adapt their strategies dynamically. For implementation, we build our environment on top of a rule-based bot [82], replacing its rule-based combat logic with the decisions made by LLM agents. To access internal game states, we utilize the Darkest Dungeon Save Editor [83]. Since the game does not support complete control via keyboard input, we employ the Xbox 360 controller emulator to inject actions from the LLM agent. We evaluate the agent’s performance during the first embarkation mission after the tutorial, which we designate as our default evaluation setting. For consistency, we fix the party roster to include the Plague Doctor, Vestal, Highwayman, and Crusader, in that order, and equip the inventory with an additional 8 ‘Food’ and 8 ‘Torches’. To ensure reproducibility, we provide a save file with this setup preconfigured. During the mission, dungeon exploration is handled by rule-based logic, while all combat decisions are delegated to the LLM agent.
@@ H.3 Evaluation Metric for Darkest Dungeon
$\text{Score}=\begin{cases}40\cdot\left(\frac{\text{\# combats cleared}}{\text{\# total combats}}\right)+30\cdot\left(\frac{\text{\# heroes survived}}{4}\right)+30\cdot\left(1-\frac{\text{total stress}}{800}\right),&\text{if stage is cleared}\\ 40\cdot\left(\frac{\text{\# combats cleared}}{\text{\# total combats}}\right),&\text{otherwise}\end{cases}$
@@ H.5 Result for Darkest Dungeon
Llama-3.2-1B 0.0$\pm$0.0 11.5
Llama-3.2-3B 47.5$\pm$39.2 9
Qwen-2.5-3B 44.8$\pm$22.2 10
Qwen-2.5-7B 88.8$\pm$2.0 6
Minitron-4B 0.0$\pm$0 11.5
Minitron-8B 63.8$\pm$30.4 8
GPT-4o-mini 81.3$\pm$5.8 7
GPT-4o 93.4$\pm$1.5 2
o3-mini 89.0$\pm$2.1 5
Gemini-2.5-pro 93.7$\pm$1.6 1
Claude-3.7 89.9$\pm$2.5 4
Deepseek-R1 91.7$\pm$1.1 3
@@
Llama-3B Zero-shot 47.5$\pm$39.2 7
Reflection 47.3$\pm$39.0 8
Planning 56.3$\pm$23.6 6
Ref-Plan 57.0$\pm$31.6 5
GPT-4o Zero-shot 93.4$\pm$1.5 1
Reflection 85.2$\pm$10.3 3
Planning 82.0$\pm$8.6 4
Ref-Plan 91.6$\pm$2.5 2
@@
GPT-4o Text 93.4$\pm$1.5 2
Image 92.2$\pm$3.0 3.5
Gemini Text 93.7$\pm$1.6 1
Image 92.2$\pm$1.8 3.5
Claude Text 89.9$\pm$2.5 6
Image 90.1$\pm$5.7 5
@@ I.1 Game Description for Minecraft
Environment. Minecraft [59] is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. This environment is based on the Mineflayer JavaScript API [60]. Using Mineflayer, the agent can control a Minecraft bot through high-level JavaScript commands. We use Minecraft version 1.19 for compatibility, and to ensure consistent evaluation, we fix the world seed to 42 and initialize the bot at coordinates (604, 100, -823). The bot starts in survival mode with no items and progressively crafts a target item using its in-game observations and JavaScript-based actions generated by the LLM. We select 8 target items with varying levels of crafting difficulty: ‘crafting table’, ‘stone pickaxe’, ‘furnace’, ‘bucket’, ‘golden sword’, ‘diamond pickaxe’, ‘enchanting table’, and ‘nether portal’.
@@
Action Space. The action space consists of JavaScript code that interfaces with the Mineflayer API. Following Voyager [20], we expose the following set of control primitives to guide the LLM in generating valid and effective code actions for the bot using in-context learning.

### v2

@@ D.5 Result for Super Mario
[…]
Minitron-8B $3±\pm1$2.8 6
[…]
@@
[…]
@@
[…]
@@ E.1 Game Description for Ace Attorney
Environment. Ace Attorney (Capcom, 2001) is a courtroom adventure game where players act as defense attorneys, gather evidence, and cross-examine witnesses. We target the first episode of Phoenix Wright: Ace Attorney Trilogy on Steam (see Figure 12). We define four subtasks: one three-question multiple-choice quiz where the player selects the correct answer, and three cross-examination tasks where the player presses witnesses for more details or presents evidence to expose contradictions. We use Harmony (Pardeike, 2025) with a BepInEx plugin (Contributors, 2025a) to hook the game’s source code at launch, capturing states such as dialogue text, arrow-button visibility, keyboard inputs, and Court Record entries. The hooks save states as .txt or .json files and monitor a command .txt file for inputs, injecting them into the game in real time.
Observation-to-Text Conversion. All states remain in text form and require only minimal post‐processing. We map speaker indices to character names (e.g., index ‘2’ $\rightarrow$ ‘Phoenix Wright’) using a predefined mapping and replace original names with arbitrary aliases to prevent contamination (e.g., ‘Phoenix Wright’ $\rightarrow$ ‘Alias’). We also convert Court Records and multiple‐choice candidate options—originally stored in .json—into continuous descriptive text when they exist.
@@ E.3 Evaluation Metric for Ace Attorney
Each subtask begins and ends at fixed points, with screenshots at both start and end and the preceding conversation history provided, and yields a reward $r_{i}$ and a step count $t_{i}$. In the multiple‐choice task (MC), $r_{i}$ is the number of correct answers out of three, while in cross‐examination tasks CE1, CE2, and CE3, $r_{i}$ is 1 for a pass and 0 for a fail; any failure incurs a maximum step count of 50. We normalize each reward and step count against fixed benchmarks $\bar{r}_{i}$ and $\bar{t}_{i}$, namely $(3,25)$ for MC, $(1,11)$ for CE1, $(1,3)$ for CE2, and $(1,4)$ for CE3, by computing
@@ E.5 Result for Ace Attorney
[…]
@@
[…]
@@
[…]
Both $\pm17.3$ 3
@@ F.1 Game Description for Her Story
Environment. Her Story (Barlow, 2015) is an interactive adventure game where players explore police interview clips to uncover a hidden truth. The player begins the game by accessing an old desktop interface, where a program called L.O.G.I.C. Database is open. The player can enter keywords into the database to retrieve up to five video clips whose transcripts contain the searched word. By watching these clips and gathering clues, the player repeatedly formulates new queries to reconstruct the underlying story. To interface the game with our code, we use Harmony (Pardeike, 2025) together with Unity Doorstop (Contributors, 2025b) to log internal game states to a .txt file. Specifically, each line in the file is a JSON object representing a snapshot of the game’s state at a given moment. Each object contains an event type and its associated metadata. The following examples illustrate key elements of the logged game state. For clarity, we omit auxiliary metadata that are present in the actual logs but are only used for debugging or UI-related purposes, such as video IDs, screen resolution, and UI element positions.
  • Load title screen: {”status”: ”title”}
  • Load L.O.G.I.C. Database: {”status”: ”start_game”}
  • Query: {”status”: ”query”, ”keyword”: keyword}
  • Get query result: {”status”: ”query_result”, ”num_total”: number of clips containing the keyword, ”num_visible”: number of clips shown, ”video”: list of {”new”: 1 if not viewed, otherwise 0, ”session”: recording date, ”outfit”: visual description of the thumbnail}}
  • Open video panel: {”status”: ”open_detail”}
  • Close video panel: {”status”: ”close_detail”}
  • Play video: {”status”: ”play_video”, ”script”: transcript of the video}
  • Close video: {”status”: ”close_video”}
@@ F.3 Evaluation Metric for Her Story
To uncover the truth behind the case, the player actively explores the video archive by issuing queries and watching clips. We evaluate this behavior by counting how many unique clips the player has viewed from the full archive of 271 clips. This aligns with the original game design, where unlocking achievements and reaching the ending depend on the number of clips watched. Specifically, we define the score as:
$\text{Score}=\left(\frac{\text{Number of distinct video clips viewed}}{271}\right)\times 100.$
@@ F.5 Result for Her Story
Table 21 summarizes the gameplay scores of different LLMs on Her Story. Commercial LLMs outperform open-source models in this task. Gemini-2.5-pro achieves the highest score of 67.3, followed by Deepseek-R1 (66.9), o3-mini (66.3), GPT-4o (64.0), Claude-3.7-sonnet (62.6), and GPT-4o-mini (21.1). The low score of GPT-4o-mini is due to its repeated search of keywords such as alibi, evidence, and witness, even after those queries have already been issued. This redundancy reduces its ability to discover new clips. Among open-source LLMs, all models score below 10. Most exhibited similar failure patterns: they fail to play unseen clips, repeatedly issue the same keywords, or search using entire sentences (e.g., for the name ”Simon”, for keywords related to Luna’s) rather than meaningful single words.
@@
Llama-3.2-1B $\pm1.2$ 11
[…]
@@
[…]
@@
[…]
Both $6±\pm3.$1 4
@@ G.1 Game Description for Pokémon Red
[…]
@@ G.3 Evaluation Metric for Pokémon Red
$\text{Score}=\big(\frac{\text{Number of flags achieved}}{12}\big)\times 100.$
@@ G.5 Result for Pokémon Red
Llama-3.2-1B $\pm0.0$ 8.5
Llama-3.2-3B $\pm0.0$ 8.5
Qwen-2.5-3B $\pm0.0$ 8.5
Qwen-2.5-7B $\pm0.0$ 8.5
Minitron-4B $\pm0.0$ 8.5
Minitron-8B $\pm0.0$ 8.5
GPT-4o-mini $\pm0.0$ 8.5
[…]
o3-mini $\pm0.0$ 8.5
[…]
@@
Llama-3B Zero-shot $\pm0.0$ 6.5
Reflection $\pm0.0$ 6.5
Planning $\pm0.0$ 6.5
Ref-Plan $\pm0.0$ 6.5
[…]
@@
[…]
@@ H.1 Game Description for Darkest Dungeon
Environment. Darkest Dungeon is a turn-based roguelike role-playing game where the player manages a roster of heroes as they explore procedurally generated dungeons filled with monsters, traps, and treasures  (Red Hook Studios, 2016). Each hero has unique abilities and a stress level that influences their behavior during combat and exploration. Stress accumulates through continued exploration and battle, and heroes who reach high stress thresholds may develop afflictions that hinder or occasionally enhance their performance. The game emphasizes tactical positioning, turn-order strategy, and long-term roster management. Elements of randomness, such as attack accuracy, critical hits, and affliction outcomes, introduce uncertainty and require players to adapt their strategies dynamically. For implementation, we build our environment on top of a rule-based bot (Kgleken, 2023), replacing its rule-based combat logic with the decisions made by LLM agents. To access internal game states, we utilize the Darkest Dungeon Save Editor (Robojumper, 2023). Since the game does not support complete control via keyboard input, we employ the Xbox 360 controller emulator to inject actions from the LLM agent. We evaluate the agent’s performance during the first embarkation mission after the tutorial, which we designate as our default evaluation setting. For consistency, we fix the party roster to include the Plague Doctor, Vestal, Highwayman, and Crusader, in that order, and equip the inventory with an additional 8 ‘Food’ and 8 ‘Torches’. To ensure reproducibility, we provide a save file with this setup preconfigured. During the mission, dungeon exploration is handled by rule-based logic, while all combat decisions are delegated to the LLM agent.
@@ H.3 Evaluation Metric for Darkest Dungeon
[…]
@@ H.5 Result for Darkest Dungeon
Llama-3.2-1B $\pm0.0$ 11.5
[…]
@@
[…]
@@
[…]
@@ I.1 Game Description for Minecraft
Environment. Minecraft (Mojang Studios, 2011) is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. This environment is based on the Mineflayer JavaScript API (contributors, 2013). Using Mineflayer, the agent can control a Minecraft bot through high-level JavaScript commands. We use Minecraft version 1.19 for compatibility, and to ensure consistent evaluation, we fix the world seed to 42 and initialize the bot at coordinates (604, 100, -823). The bot starts in survival mode with no items and progressively crafts a target item using its in-game observations and JavaScript-based actions generated by the LLM. We select 8 target items with varying levels of crafting difficulty: ‘crafting table’, ‘stone pickaxe’, ‘furnace’, ‘bucket’, ‘golden sword’, ‘diamond pickaxe’, ‘enchanting table’, and ‘nether portal’.
@@
Action Space. The action space consists of JavaScript code that interfaces with the Mineflayer API. Following Voyager (Wang et al., 2023), we expose the following set of control primitives to guide the LLM in generating valid and effective code actions for the bot using in-context learning.